OpenStudy (anonymous):

1/6a=1 1/3 , solve and check

OpenStudy (saifoo.khan):

\[\frac16a = 1\frac13\]Like that?

OpenStudy (anonymous):

yepp

OpenStudy (saifoo.khan):

\[\frac16a = \frac43\]Cross multiply. \[a = \frac{4 \times 6}{3}\] \[a = 4 \times 2\] a= 8

OpenStudy (anonymous):

and how would i check ?

OpenStudy (phi):

replace a with 8: 8/6 =?

OpenStudy (anonymous):

okay